Prev Next

SW&IT Meeting Minutes 2018-09-10

Attendees: David, Igor, Kellee, Matthieu, Petr, Stephen, Xianju, Yang

Outlook

News from management

  • s/w recruitment
  • finance committee cybersecurity report
  • Peter G moving in 508e

Previous Action Items

  • Vsphere (DE)
  • power plan update draft (DE)
  • Luci H/O doc (IS)
  • Luci telemetry - web-io cold heads state for vibration (IS)

Transition plan

Handover steps

  1. primary + secondary do self learning, gather information, read wiki
  2. primary organize teamviewer/zoom walk through with Kellee and secondary
  3. primary draft techtalk w. input from secondary
  4. techtalk review by Kellee
  5. primary does a team presentation

Please update this table to indicate progress.


Application(s) Primary Secondary WeekSorted ascending
OAC
AC (2)
Petr (2)
1
ALH Igor(1/2)
Matthieu (2)
1
GCS
Petr (2)
Matthieu (2)
1
AzCam
AC (2)
Petr (2)
2
DIMM
AC (2)
Petr (2)
2
Telemetry
Petr (1)
Matthieu (1)
2
Status Server Stephen (1) Matthieu 2
LBC
Yang (1)
Igor(1/2) 2
MODS Yang Igor 2
Weatherstation AC Petr 3
INDI Igor Matthieu 3
OVMS
Petr (2)
Matthieu (2)
3
Telemetry cleanup Stephen Petr 3
Telem-VIZ Yang Stephen 3
FACSUM,WEATHERSTATION AC Matthieu 4
particle monitors Yang Petr 4

Open topics

Bring any topics you want for discussion at the meeting here. Use <topic>[<name>] formatting.

  • UA time
  • Telescope time, mountain managers

This week in github

Team reports

MB

  • management OC, WPC chair
  • Handover activities
  • Science Archive digging into sciops
  • s/w interview
  • cyber security report for finance committee
  • nightwriter discussion with Shane Walsh

next:
  • IT to github issues
  • TMT ESW review
  • jumbo / Soul

YZ

  • Working on transition to Kellee's LBC responsibility.
    • Met with Kellee, Igor, and Olga.
    • Editing notes with Igor.
    • Will meet with Kellee again sometime.
    • LBC is complicated so takes longer time (will start MODS learning afterwards).
  • Working on collimation software.
  • Continue working on documenting pointing and collimation procedures.
  • Continue working on Python package and modules.
  • Need to start gather more info on Al's AO project.

PK

  • Chasing two TCS problems - left WFS not converging and LBC NS tracking
  • Upgraded GCSGUI, working on more updates (replacing QLabel with proper QGraphicsView,..)
  • AZCam, OAC and DIMM HO meetings
  • Edited OVMS kernel update wiki

next:
  • Further chasing (and hopfully solving) TCS problems
  • Telemetry and Weatherstation HO
  • GCS, OVMS presentation

KS

Last week
  • Weatherstation issue - not reading wind data. LeRoy found the media converters in a bad state and one fiber broken. Rear is still not reading.
  • Transition meetings - ALH, OAC, AzCam, LBC, DIMM.
  • Wiki updates for OAC telemetry and new build dependencies.
  • LBC blue tech controller reading bad voltages - a little delving into the code to see how it's checking.

This week
  • Continue handover meetings as they come up.
  • Walk through LBC log with Yang/Igor.

SH

Last week:
  • Jumbo frames switch configuration, CRA configuration, LBTI computer changes.
  • clean-up of rm507 computers.
  • Cindy e-mail change (Google makes this really easy)
  • obs1 running VNC for mirroring
  • jet issues on Thursday

This week:
  • Statserv handover
  • Jumbo frames meeting and network changes continued.
  • Mountain on Thursday to make sure SOUL jumbo frames ready to go.
  • Changes to networking for MGIO in utility building.

IS

Last Week:
  • HandOver
    • ALH:
      • Going through documents and code. Had meeting with Keelee and Mathieue
    • LBC:
      • Went through some highlevel documentation and view some of the code, part because of handover and part because of #7380
  • PIT
    • Revisit PIT table and send a test release to science operations.
  • OT/LBC #7380
    • Possible issue with proper motion units/amounts in LBC scripts. As part of the investigation had to generate a spreadsheet with all the observations done since proper motion was added to the OT and grab the LBC XML generated. Also started to look a little bit into the LBC code to understand how proper motion is handle.

This Week:
  • LUCI development of AFC RTD
  • Handover
  • PIT point release.

DE

Last week:
  • Followup on DIN switches. Should ship today, Perle upgraded us to overnight shipping
  • VMWare upgrade consult
    • Direct path to 6.5 U2
    • Latest is 6.7, but I don't see a reason to double-update right now
  • Handholding Ketiv with Vault credentials
    • Vault looks healthy
  • Changed setting for Groups access at Christian's request to allow external committee members to use Groups
  • Peter Gray setup tweaks
    • Should have his LBT computer this week
  • Meeting with UPS vendor moved to Tuesday, should have good info then
This week:
  • Go through VMWare upgrade docs to double-check for gotchas, though I think we are okay
  • Set up Peter's computer/office
  • Meeting with Tripp-Lite re: 3-phase UPS
  • OAC telemetry sync script
Topic revision: r20 - 10 Dec 2018, UnknownUser
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback